# Ginkeepaspectratio=true,clip=true,draft=false

### v1

@@ 6.3.1 Networking
We now perform an experiment to investigate the impact of the cost of microkernel IPC operations. Mi et al. [2019] compare IPC costs between seL4, Fiasco.OC and Google’s Zircon microkernel on an x86 platform. They find intra-core fastpath round-trip IPC latencies for the three kernels to be 986, 2717 and 8157 cycles respectively. So we simulate those kernels by adding $(2717-986)/2=865$ (”Fiasco”) and $(8157-986)/2=3585$ (“Zircon”) respectively to each seL4 system call (by executing a tight loop while monitoring the cycle counter). We do not claim that this is in any way precise, but it should give a rough estimate of how the same setup will perform on those kernels.
Figure 10 shows the results. We see that the seL4-based LionsOS has lower CPU usage and achieves higher throughput than the Fiasco simulation. Simulated Zircon only achieves about 2/3 of the seL4 throughput.
Figure 11: Ping latencies on system with three prioritised clients as a function of load on the medium-priority client on Arm.
Virtualiser policy
Figure 11 shows an experiment with three clients, to which the TxVirt assigns different policies (high, medium, low). We vary the load on the medium-priority client and impose no load on the others, other than pinging. The graph shows ping latencies of the three clients as a function of load on the medium-priority client. The priorities are reflected in the ping times being fastest on the high-priority, and slowest on the low-priority client.
Figure 12: Two client system demonstrating virtualiser swapping and throughput limiting.
@@
We investigate swapping policies at run time. To demonstrate this we configure a system with a TxVirt that monitors bandwidth, plus a Swapper that can reload the TxVirt with a different executable, a bandwidth limiter. When a preset throughput threshold is exceeded, the Virt signals the Swapper, which the loads the new program into the Virt. (This setup uses experimental Microkit features.)
Figure 12 shows the system running on the Arm platform. Client 0 has a threshold of 500 Mb/s, and is throttled to 500 Mb/s. Client 1 has a threshold of 200 Mb/s and also throttled to 200 Mb/s. The swap happens at the point where the clients reach 200 Mb/s. The graph clearly shows the bandwidth limitation working as intended. We measure the time to perform the switch to be 17 $\mu$s.
6.3.2 Block device
Figure 13: Random read bandwidth on LionsOS and Linux.
We compare our block device driver and virtualiser with raw reads from Linux on a Samsung SDXC card, see Figure 13. Performance is essentially identical on the two platforms, and limited by the card’s performance. Linux does a slightly better job with small block sizes on a random read workload, due to LionsOS presently handling the device sub-optimally. However, LionsOS per-block processing cost is vastly lower than that of Linux.
@@ 7 Conclusions
To summarise our experience with building and evaluating LionsOS:
  • It is possible to build an OS based on strict separation of concerns, without sacrificing performance. In fact, LionsOS significantly outperforms Linux (and by implication any other modular OSes for which performance data are available).
@@
  • The simplicity of design and implementation makes us confident that the verification aim is achievable. The from-scratch implemented components that make up the system’s TCB are all far simpler than other systems verified with automated provers, and similar to the SMT-verified the Microkit.
  • The adaptability aim also seems in reach: use-case specific policy modules are simple and easy to swap. Good tooling will be required to help developers design and tailor a system for their use case.
  • All this is achieved without changing a single line of code in seL4. In particular, we see no need for shortcuts and compromises such as migrating critical functionality into the kernel, or co-locating services, as Chen et al. [2024] found necessary.

### v2

@@ 6.3.1 Networking
We conduct experiments to evaluate the impact of context-switching/IPC costs. For we first use IPC costs of the seL4, Fiasco.OC (now called L4Re) and Google’s Zircon microkernels as measured by Mi et al. [2019]. They found intra-core fastpath round-trip IPC latencies for the three kernels to be 986, 2717 and 8157 cycles respectively. So we simulate Fiasco and Zircon kernels by adding $(2717-986)/2=865$ (”Fiasco”) and $(8157-986)/2=3585$ (“Zircon”) respectively to each seL4 system call (by executing a tight loop while monitoring the cycle counter). We do not claim that this is in any way precise, but it should give a rough estimate of how the same setup will perform on those kernels. We also estimate performance with “free” IPC by subtracting $986/2=493$ cycles from every system call.
5(d) shows that while higher IPC costs lead to higher CPU load, all three kernels manage to handle the applied load (barely in the case of “Zircon”). Making IPC free has minimal impact on LionsOS performance.
6(a) shows a second experiment, where we compare our standard unicore configuration (with the RxCopy component, “Copy” in the graph) to a configuration without Copy (“none”) or with a “Null” component that simply forwards packets without any processing. The difference between Null and Copy is the cost of copying the packet, while “none” vs. “Null” shows the context-switch overhead, as does comparing “Copy” with “Copy+Null”. As the figure shows, the cost of an extra context switch is almost invisible.
For the “free” IPC case we subtract an estimate of the kernel costs, simulating an ideal case where module switches are totally free (in reality there would still be function-calls). The difference is highest at low load (natural batching reduces the context-switch rate at high loads, up to two dozen packets are handled per invocation). The difference between this over-idealised case and the normal configuration is still less than the difference between the latter and Linux.
(a) Sequential read.
(b) Random read.
(c) Sequential write.
(d) Random write.
(e) Reading multiple files.
Figure 7: I/O operation bandwidth on LionsOS and Linux. Solid lines are throughput, dashed lines are CPU utilisation.
Virtualiser
To evaluate the scalability of sharing the NIC between multiple clients, we vary the number of clients served by the Virts from one to 32. 6(b) shows CPU usage as a function of applied load, when only one client has traffic directed to it. There is only a slight dependence of CPU use on the number of clients, resulting from the TxVirt having to poll all queues.
6(c) shows the same scenario, except that the load is now evenly distributed between the clients. The performance impact of client number is higher than in the previous case, presumably a result of the higher cache footprint of the multiple clients. In embedded systems, the number of network clients will be small, so this is an encouraging result.
@@
We investigate swapping policies at run time. To demonstrate this we configure a system with a TxVirt that monitors bandwidth, plus a Swapper that can reload the TxVirt with a different executable, a bandwidth limiter, which is on standby. When a preset throughput threshold is exceeded, the Virt signals the Swapper, which then loads the new program into the Virt.
6(d) shows the system running on the Arm platform. We have two clients running in the system. Client 0 has a constant applied load of 10 Mb/s, while Client 1 has an increasing load. When this reaches a threshold of 500 Mb/s at about 1.7 s, Client 1 is throttled to 200 Mb/s through a virtualiser swap. The graph clearly shows the bandwidth limitation working as intended, Client 0 only experiences an increased variance in its latency. It takes 17 $\mu$s to perform the switch, which is so short that Client 0 does not experience dropped packets or even a temporary latency spike.
6.3.2 Storage
Raw I/O
We evaluate storage-driver and Virt performance on the Odroid-C4 board using a Sandisk SDXC card. To minimise the effect of the device’s flash management, we use a different area of the card for each benchmark run, separated by $>$16 MiB. We power-cycle the card between runs, to clear any in-card cache.
As shown in Figure 7, throughput of LionsOS and Linux are similar, and limited by the card’s performance. However, for small block sizes, where software overheads matter, LionsOS achieves this with a fraction of the CPU used by Linux.
File system
The performance of reading a number of small files can matter in some systems [Chen et al., 2024]. We therefore conduct a benchmark where we open-read-close a number of files on a FAT32 filesystem using blocking I/O on both LionsOS and Linux running on the Odroid-C4 board.
The benchmark reads 4 files of size 512 bytes, all in the same directory two levels from the root directory. 7(e) shows the results compared to Linux. The time for reading each file is the same in both systems. Linux takes a long time to open the first file, but this is amortised when accessing multiple files – showing the cost and benefits of directory entry caching. LionsOS presently does no directory caching, and is slower when reading more than five files. LionsOS takes far longer than Linux to close files, a result of all resource management being done synchronously, while Linux defers clean-up.
7 Discussion
Our evaluation shows that LionsOS is able to achieve excellent performance in networking, comprehensively outperforming Linux in achieved throughput, latency and processing costs. This performance is achieved despite each packet being handed between several modules, each time requiring a context switch. Our analysis of the impact of context-switch/IPC costs (Section 6.3.1) shows that these only have a minor impact on overall performance, and the additional benefit from a hypothetical zero-cost IPC cost is small compared to the already demonstrated performance advantage over Linux. This contradicts claims in the literature that IPC cost is still a bottleneck in microkernels [Chen et al., 2024].
Storage I/O performance shows the same picture: Raw I/O bandwidth equals Linux’ (or is slightly better) with a significantly reduced processing cost for small block sizes (where software overheads dominate).
The main reason for the high performance is that each component is very simple, doing just one thing. There is no need for any interface layers for adapting between multiple implementations of the same functionality; most of the components are simple straight-line code with no callouts via function pointers. This is possible thanks to the simplicity afforded by keeping policy implementations specialised to the use case.
The file-read benchmarks indicate that directory caching would help where many files are being accessed in the same directory. This is an example where the best policy depends on the use case, and a poor policy choice may be detrimental. As such it is an example where a use-case-specific caching policy should be implemented as part of the directory management.
The experiments with swapping a component at run time (Section 6.3.1) show that dynamically changing between pre-configured policies can be done very quickly (17 $\mu$s on our Arm platform).
@@ 8 Conclusions
Summarising our experience with building and evaluating LionsOS:
  • It is possible to build an OS based on strict separation of concerns, without sacrificing performance.
@@
  • The simplicity of design and implementation makes us confident that the verification aim is achievable. The from-scratch implemented components that make up the system’s TCB are all far simpler than other systems, in fact, a high-performance NIC driver has already been verified using automated proof techniques [Zhao et al., 2025].
  • Contrary to other claims [Chen et al., 2024] we find that IPC cost matters little in our design.